OpenStudy (anonymous):

Sharmi and Reese each deposit money into a savings account. Sharmi deposits $300 less than Reese. After the same amount of time, Reese has earned $180 in interest at a rate of 3% per year, and Sharmi has earned $270 in interest at a rate of 6% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (anonymous):

ok.....let reese deposit say $x , then sharmi would have deposited $(x -300)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Lets calculate the interest for both separately...Is for Sharmi and Ir for reese

OpenStudy (anonymous):

note that the time value t is same for both of them

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lets start with reese, the formula is I = PRT/100

OpenStudy (anonymous):

so we have 180 = (x)(t)(3)/100

OpenStudy (anonymous):

180 = .03 t x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

similarly 270 = .06t (x - 300)

OpenStudy (anonymous):

so we have 2 equations in 2 unknown variables

OpenStudy (anonymous):

okay, so solve for x?

OpenStudy (anonymous):

yepp...am solving....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

from the first equation, x = 180 /.03t

OpenStudy (anonymous):

substitute this value of x in the second equation and we get 270 = 360 - 18t

OpenStudy (anonymous):

that is t = 5

OpenStudy (anonymous):

okay, so 270=360-18t -90=-18t t=5

OpenStudy (anonymous):

substitute the value of t in the first equation and we get 180 = .03 (5) x

OpenStudy (anonymous):

x = 180/.15

OpenStudy (anonymous):

x is 1200

OpenStudy (anonymous):

Reese initially deposited $324, Sharmi initially deposited $24. Reese initially deposited $900, Sharmi initially deposited $600. Reese initially deposited $1,200, Sharmi initially deposited $900. Reese initially deposited $1,620, Sharmi initially deposited $1,300.

OpenStudy (anonymous):

those are the answer choices, so the answer is the third one?

OpenStudy (anonymous):

i guess its the third one

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thats what i arrived at x is 1200 and x - 300 is 1200 - 300 that is 900

OpenStudy (anonymous):

I got x = 1200... so, yes, 1200 and 900...